# Query Rewriting via LLMs

### v2

@@ 7.1.2 Execution Time
Thus far, we had considered optimizer-estimated execution costs. We now move to wall-clock runtimes for query executions – Figure 8 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA. We observe, again gratifyingly, that there are indeed several queries where substantial time benefits are achieved by the rewrites, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a huge factor of 10000! Second, in almost all cases, LITHE outperforms or matches SOTA, including as mentioned above, the queries where SOTA’s optimizer costs were better. There remain only a few instances (Q9, Q47) where SOTA is marginally better.
From a modeling perspective, we see that the well-documented gap between optimizer predictions and actual run-times is prevalent in the rewrite space as well. On the one hand, there is Q45 where the projected speedup of $10^{2}$ increases to a huge $10^{4}$ at runtime. On the other hand, the $10^{5}$ speedup for Q41 decreases to $10^{3}$. But for SOTA, the reductions can be severe – a striking case in point is Q57, where SOTA actually causes regression despite a speedup projection of close to 100. However, the good news is that with LITHE, although the runtime speedups did not always match the projections, we did not encounter any regressions among the CPR rewrites. Overall, LITHE produces more robust rewrites resulting in a high TSGM of 18.4, whereas SOTA provides a TSGM of 6.
@@ 7.1.3 Reasoning
As a confirmatory exercise, we compared the LLM-generated explanation output by LITHE with our own manual analysis of the query plans generated for the original and rewritten queries. For example, consider Q45 – the provided explanation correctly lists the main reasons for the huge $10^{4}$ speedup as “the use of CTEs”, “Reduction in data volume early on” through pre-filtering based on join predicates, etc. Overall, we found that the explanations provided by LITHE matched our manual analysis of the plans, indicating that model-based reasoning is well aligned with human-backed reasoning in these scenarios.
@@ 7.3.1 Transformation Time
The average processing time per CPR query is shown in Table 6, and we see that they do take a few minutes. However, this investment may be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads. For instance, with Q11, the original query took nearly two hours to complete, whereas the LITHE rewrite executed in just 3 minutes. Further, many applications tend to use a set of canned queries which are run thousands of times. Thus, even a large one-time investment can be easily recovered over repeat executions of such queries.
@@
Notwithstanding the above, we also observe that LITHE is considerably slower than SOTA in producing rewrites. A drill-down shows that the lion’s share of the time is taken by the initial prompt ensemble and the final MCTS module. We explore techniques to improve their efficiency in the Section B.
@@ 7.4 Commercial DBMS
LITHE 7 8 3.6 4.8 1.5 2.4
SOTA 3 6 2.1 3.5 1.3 1.7
The performance of LITHE and SOTA on these two systems is shown in Table 7, with LITHE continuing to do better than SOTA. Despite the apparent lack of optimization headroom, LITHE still produces 7 and 8 CPRs resulting in a healthy CSGM of 3.6 and 4.8, respectively. Further, the TSGM provided by these rewrites are a useful 1.5 and 2.4, respectively. However, unlike the no-regression outcome seen with PostgreSQL, LITHE generates 3 CPRs on OptA, and 2 on OptB, which turn out to be regressions. So, although these commercial optimizers may have greater sophistication, they also suffer significant brittleness with regard to estimation accuracy.
Interestingly, one regression on OptB was caused by the LLM bringing in, via a basic prompt, an unrestricted version of Rule 6, which is prone to generating poorly-estimated plans (see Section 8.3). Despite such occasional errors, LITHE provides good overall improvement in runtimes as indicated by the TSGM values.
We also performed a preliminary study of LITHE on an internal real-world benchmark, whose data and query characteristics are more realistic than TPC-DS. Even on this hard benchmark, LITHE produced 8 CPRs (out of 45 queries tested) with a CSGM of 2 and a TSGM of 3.3.
@@ 7.5.1 Masked Database
Approach # CPR CSGM
TPC-DS Masked TPC-DS Masked
SOTA 15 12 6.1 5.3
LITHE 26 24 11.5 10.5
@@ 7.5.2 LITHE on LLaMA
Table 9: LITHE Rewrite Performance with LLaMA.
@@
LLaMA without MCTS 17 5.6 8.2
LLaMA with MCTS 21 8.3 14.7
@@ 8.2 Rewrite Migration to Optimizer
8.3 Revisiting Optimizer Plan Costing
As highlighted in Section 7, there were instances of substantive differences between the promised speedup and that delivered at run-time. In fact, to the extent that speedups could even turn out to be regressions! This is due to the brittleness of optimizer plan costing in the new spaces explored by the LLM. As a case in point, during our LITHE design process, the prefiltering in Rule R6 (Table 1) had initially not been restricted to self-joins. It resulted in the number of CPRs (on PostgreSQL) being as high as 65, with an astonishing CSGM of 30.6! However, upon execution, most of these rewrites turned out to be regressions, which led to our inclusion of the restriction. But note that we are artificially modulating the rewrite, rather than fixing the plan costing module, which is the principled solution. In sum, while plan cost modeling has been a long-standing area of research, there is now even more reason given the new rewrite spaces to study this topic further – for instance, the operator cost model could be extended using calibration techniques similar to those advocated in [40], while the operator cardinality model could be improved with attention-based techniques [23].
@@ 8.5 Scope of Semantic Equivalence Tools
As seen in the experiments section, the coverage of logic-based query equivalence testing on industrial-strength queries is rather limited. On the other hand, while it is highly likely that the statistics-verified rewrites are valid, it still requires the DBA to make a final call on the correctness. This limitation restricts the use of LITHE in a fully automated scenario, i.e., as a direct preprocessor to the query engine. Therefore, a key challenge is to improve logic-based coverage.
@@ 9 Related Work
Rule-based SQL rewriting. Most of the recent work on SQL query rewriting is rule-based [45, 38, 8, 39, 10, 28]. For instance, WeTune [38] employs a rule generator to enumerate a set (up to a maximum size) of logically valid plans for a given query to create new rewrite rules, and uses an SMT solver to prove the correctness of the generated rules. While this approach can generate a large set of new rewrite rules, it often fails in coming up with transformation rules for complex queries due to the computational overheads of verifying rule correctness. As such, it is unable to rewrite any of the TPC-DS queries [16, 25].
Learned Rewrite [45] uses a set of existing rewrite rules from Calcite [9] and aims to learn the optimal subset of rules along with the order in which they must be applied. Given that the search space for possible rewrites is exponential in the number of applicable rules, an MCTS scheme is used to efficiently navigate this space and find the rewritten query with maximum cost reduction.
@@
LLM-based rewriting. GenRewrite [25] is the first LLM-based approach that tries to use the LLM itself for end-to-end query rewriting. Instead of using predefined rules from Calcite [9], they employ the LLM to create Natural Language Rewrite Rules (NLR2s) to be used as hints, and perform several iterations of prompting to get the rewritten query. They show that LLMs can outperform rule-based approaches due to their ability to understand contexts, and demonstrate a significant improvement in query rewriting compared to prior methods.
@@ 10 Conclusions and Future Work
Appendix A Rewrite Quality (Other Benchmarks)
@@ Appendix B Overheads Reduction for LITHE
Table 11 compares the performance of LITHE with and without the classifier. The time overheads do visibly go down by about 52 percent, and the tokens by about 13 percent. However, there is a price to be paid – the CPR is reduced to 24 and the CSGM comes down to 8.5. In our future work, we plan to look into whether a better tradeoff could be achieved between quality and overheads.
@@ Appendix C Query Equivalence Testing via Sampling
If the original query is not within the scope of the logic-based QED tool, we alternatively use a sampling-based approach to test equivalence. The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not a sufficient condition. That is, there are no false negatives, but there can be false positives. This is because the sampled database may not cover all the predicates present in the query. This can cause two types of problems: First, it is possible for two different queries to return the same result. This can happen when, for example, the entirety of the sampled data satisfies a predicate of one query, while the same predicate is not present in the other. Second, if the underlying sample does not satisfy any of the predicates in either query, then an empty result will be returned by both queries. This again does not imply that the queries are equivalent.
@@
    To differentiate between inner and outer joins, we insert into the sample database, rows with appropriate values on the relevant columns such that they produce NULL values for outer joins.
@@
    Given a pair of queries to test for equivalence, we adjust the constants in the filter predicates to reduce the chances of an empty result. We make use of the rows in the sampled data for this purpose. For example, say an equality predicate is present in the query and the associated constant is absent in the sampled database. We then replace the query constant with a value already present in the sample. Similarly, the constants for other comparison operators are adjusted based on the ranges of the corresponding columns in the sampled database. Note that these modifications only change the selectivity of the query, but not its semantics.
@@ Appendix D Detailed MCTS Algorithm
11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n,a)$

### v3

@@ 7.1.2 Execution Time
Thus far, we had considered optimizer-estimated execution costs. We now move to wall-clock runtimes for query executions – Figure 8 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA on a cold-cache environment. We observe, again gratifyingly, that there are indeed several queries where substantial time benefits are achieved by the rewrites, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a large factor of 700! Second, in all cases, LITHE outperforms or matches SOTA, including as mentioned above, the queries where SOTA’s optimizer costs were better.
From a modeling perspective, we see that the well-documented gap between optimizer predictions and actual run-times is prevalent in the rewrite space as well. On the one hand, there is Q45 where the projected speedup of 40X increases to 700X at runtime. When run in a warm-cache environment, this speedup increases to a huge $10^{4}$ (see Section A). On the other hand, the $10^{5}$ speedup for Q41 decreases to a still reasonable 200X. But for SOTA, the reductions can be severe – a striking case in point is Q57, where SOTA actually causes regression despite a speedup projection of close to 100. However, the good news is that with LITHE, although the runtime speedups did not always match the projections, we did not encounter any regressions among the CPR rewrites. Overall, LITHE produces more robust rewrites resulting in a high TSGM of 13.2, whereas SOTA provides a TSGM of 4.9.
@@ 7.1.3 Reasoning
As a confirmatory exercise, we compared the LLM-generated explanation output by LITHE with our own manual analysis of the query plans generated for the original and rewritten queries. For example, consider Q45 – the provided explanation correctly lists the main reasons for the 700X speedup as “the use of CTEs”, “Reduction in data volume early on” through pre-filtering based on join predicates, etc. Overall, we found that the explanations provided by LITHE matched our manual analysis of the plans, indicating that model-based reasoning is well aligned with human-backed reasoning in these scenarios.
@@ 7.3.1 Transformation Time
The average processing time per CPR query is shown in Table 6, and we see that they do take a few minutes. However, this investment may be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads. For instance, with Q11, the original query took nearly an hour to complete, whereas the LITHE rewrite executed in just 3 minutes. Further, many applications tend to use a set of canned queries which are run thousands of times. Thus, even a large one-time investment can be easily recovered over repeat executions of such queries.
@@
Notwithstanding the above, we also observe that LITHE is considerably slower than SOTA in producing rewrites. A drill-down shows that the lion’s share of the time is taken by the initial prompt ensemble and the final MCTS module. We explore techniques to improve their efficiency in Section B.
@@ 7.4 Commercial DBMS
LITHE 12 9 3.6 4.1 2.1 1.9
SOTA 3 5 1.5 1.3 1.4 1.2
The performance of LITHE and SOTA on these two systems is shown in Table 7, with LITHE continuing to do better than SOTA. Despite the apparent lack of optimization headroom, LITHE still produces 12 and 9 CPRs resulting in a healthy CSGM of 3.6 and 4.1, respectively. Further, the TSGM provided by these rewrites are a useful 2.1 and 1.9, respectively.
Interestingly, although we did not observe any regressions with PostgreSQL, a few did surface in the commercial systems. Nevertheless, our regression identification mechanisms effectively caught these brittle rewrites. As a case in point, a promising rewrite, as estimated by the optimizer, for Q23, actually takes 37 minutes to complete as compared to 18 minutes for the original query – this doubling slowdown was successfully flagged by the SEER heuristic, and the rewrite was abandoned.
We also conducted a preliminary study of LITHE on an internal real-world benchmark with more realistic data and query characteristics than TPC-DS. Even on this benchmark, LITHE produced 8 CPRs (out of 45 queries tested) with CSGM of 2 and TSGM of 3.3.
@@ 7.5.1 Masked Database
Approach # CPR CSGM TSGM
TPC-DS Masked TPC-DS Masked TPC-DS Masked
LITHE 26 24 11.5 10.5 13.2 11.8
SOTA 13 12 6.1 5.3 4.9 4.2
@@ 7.5.2 LITHE on LLaMA
Table 9: LITHE Rewrite Performance with LLaMA
@@
LLaMA without MCTS 18 5.6 6.5
LLaMA with MCTS 22 8.5 10.9
@@ 8.4 Scope of Semantic Equivalence Tools
As seen in the experiments section, logic-based query equivalence testing covers industrial-strength queries only to a limited extent. On the other hand, while it is highly likely that the statistics-verified rewrites are valid, it still requires the DBA to make a final call on the correctness. This limitation restricts the use of LITHE in a fully automated scenario, i.e., as a direct preprocessor to the query engine. Therefore, a key challenge is to improve logic-based coverage.
@@ 9 Related Work
Rule-based SQL rewriting. Most of the recent work on SQL query rewriting is rule-based [48, 42, 9, 43, 11, 31]. For instance, WeTune [42] uses a rule generator to enumerate a set (up to a maximum size) of logically valid plans for a given query to create new rewrite rules, and uses an SMT solver to prove the correctness of the generated rules. While this approach can generate a large set of new rewrite rules, it often fails in coming up with transformation rules for complex queries due to the computational overheads of verifying rule correctness. As such, it is unable to rewrite any of the TPC-DS queries [18, 28].
Learned Rewrite [48] uses existing Calcite [10] rules and aims to learn the optimal subset of rules along with the order in which they must be applied. Since the rewrite search space grows exponentially with the number of rules, it uses MCTS scheme to efficiently navigate this space and find the rewritten query with maximum cost reduction. This is in contrast to LITHE where MCTS is used to search over the output token space of an LLM in order to prevent LLM hallucinations. Specifically, our use of MCTS ensures that the LLM correctly follows the prompt instructions.
@@
LLM-based rewriting. GenRewrite [28] is the first LLM-based approach to use the LLM for end-to-end query rewriting. Instead of using predefined rules from Calcite [10], they employ the LLM to create Natural Language Rewrite Rules (NLR2s) to be used as hints, and perform several iterations of prompting to get the rewritten query. They show that LLMs can outperform rule-based approaches due to their ability to understand contexts, and demonstrate a significant improvement in query rewriting compared to prior methods.
@@ 10 Conclusions and Future Work
Appendix A Additional Experiments
A.1 Rewrite Quality (TPC-DS) Contd.
Figure 10 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA for the queries run on a warm-cache environment.
Figure 10: Execution Time Speedups in a warm-cache scenario.
As expected, the speedups often significantly increase when compared to the cold-cache scenario. For instance, the speedup for Q45 increases from 700X to over $10^{4}$. Overall, the TSGM of LITHE increases to 17.1 (from 13.2), while that of SOTA increases to 5.7 (from 4.9).
A.2 Rewrite Quality (Other Benchmarks)
@@ Appendix B Overheads Reduction for LITHE
Table 11 compares the performance of LITHE with and without the classifier. The time overheads do visibly go down by about 52 percent, and the tokens by about 13 percent. However, there is a price to be paid – the CPR is reduced to 23 and the CSGM comes down to 8.5. In our future work, we plan to look into whether a better tradeoff could be achieved between quality and overheads.
@@ Appendix C Query Equivalence Testing via Sampling
We use a sampling-based approach to quickly test equivalence in the rewrite generation stages of the pipeline. The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not a sufficient condition. That is, there are no false negatives, but there can be false positives. This is because the sampled database may not cover all the predicates present in the query. This can cause two types of problems: First, it is possible for two different queries to return the same result. This can happen when, for example, the entirety of the sampled data satisfies a predicate of one query, while the same predicate is not present in the other. Second, if the underlying sample does not satisfy any of the predicates in either query, then an empty result will be returned by both queries. This again does not imply that the queries are equivalent.
@@
    Given a pair of queries to test for equivalence, we adjust the constants in the filter predicates to reduce the chances of an empty result. We make use of the rows in the sampled data for this purpose. For example, say an equality predicate is present in the query and the associated constant is absent in the sampled database. We then replace the query constant with a value already present in the sample. Similarly, the constants for other comparison operators are adjusted based on the ranges of the corresponding columns in the sampled database.
    Note that these modifications only change the selectivity of the query, but not its semantics.
@@
    To cover predicate boundary conditions, we inject query-specific synthetic tuples into the sample—similar to the approach used by the XData mutant-killing tool [34]—in order to reduce the likelihood of non-equivalent queries producing the same results.
@@ Appendix D Detailed MCTS Algorithm
11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n_{cur},a)$